# Slender Trees and the Approximation Property

### v1

@@ 3. Strong Distributivity and Orders on Products
For better readability, we prove the result in a series of statements, showing where we apply which property. We let PP stand for the projection property, RP for the refinement property and MP for the mixing property. We check two cases:
@@
$(MP)\;\;$ $\exists p_{0},p_{1},q^{\prime}(((p^{\prime},q^{\prime})R(p^{\prime},q))\wedge((p_{0},q^{\prime})R(p^{\prime},q_{0}))\wedge((p_{1},q^{\prime})R(p^{\prime},q_{1}))$
$p_{0},p_{1}b(R)p^{\prime}$
$(p^{\prime},q^{\prime})R(p^{\prime},q)R(p,q)$
@@
$((p_{0},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{0})\wedge((p_{1},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{1}$
@@
$\exists(p_{0},q^{\prime}),y_{0}(((p_{0},q^{\prime})\Vdash\dot{x}=\check{y}_{0})\wedge((p_{0},q^{\prime})R(p,q)))$
@@
$p_{0}b(R)p$
$(RP)\;\;$ $((p_{0},q^{\prime\prime})R(p_{0},q^{\prime\prime\prime})R(p_{0},q^{\prime}))$
@@
Proof.
Let $\dot{f}$ be a $\mathbb{P}\times\mathbb{Q}$-name for a function such that some $(p,q)$ forces $\dot{f}\notin V$ and $\dot{f}\upharpoonright\check{u}\in V$ for every $u\in[V]^{<\delta}\cap V$. We will construct a winning strategy for INC in the completeness game of length $\delta$ played on $(\mathbb{Q},s(R,p))\upharpoonright q$. In any run of the game, we will construct $(p_{\gamma}^{0},p_{\gamma}^{1},y_{\gamma})_{\gamma\in\operatorname{Odd}}$ such that
@@
    $y_{\gamma}\in[V]^{<\delta}\cap V$ and the sequence $(y_{\gamma})_{\gamma\in\operatorname{Odd}}$ is $\subseteq$-increasing
@@
Assume the game has been played until some even ordinal $\gamma<\delta$. Let $y_{\gamma+1}^{\prime}:=\bigcup_{\alpha\in\gamma\cap\operatorname{Odd}}y_{\gamma}$, which has size $<\delta$. $\dot{f}\upharpoonright y_{\gamma+1}^{\prime}$ is forced to be in $V$, so we can find $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime})R(p,q_{\gamma})$ which decides $\dot{f}\upharpoonright y_{\gamma+1}^{\prime}$. By the projection property, we can find $q_{\gamma+1}^{\prime\prime}$ such that $(p,q_{\gamma+1}^{\prime\prime})R(p,q_{\gamma})$ and $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime})$, so $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ also decides $\dot{f}\upharpoonright\check{y}_{\gamma+1}^{\prime}$. Because $\dot{f}$ is forced to be outside of $V$, there is $\alpha$ such that $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ does not decide $\dot{f}(\check{\alpha})$. Let $y_{\gamma+1}:=y_{\gamma+1}^{\prime}\cup\{\alpha\}$. Then $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ does not decide $\dot{f}\upharpoonright\check{y}_{\gamma+1}$, so we find $q_{\gamma+1}^{\prime\prime\prime}$ and $p_{\gamma+1}^{0},p_{\gamma+1}^{1}$ such that $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime\prime})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ and $(p_{\gamma+1}^{0},q_{\gamma+1}^{\prime\prime\prime})$ and $(p_{\gamma+1}^{1},q_{\gamma+1}^{\prime\prime\prime})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma+1}$ differently. Lastly, use the projection property to obtain $q_{\gamma+1}$ such that $(p,q_{\gamma+1})R(p,q_{\gamma+1}^{\prime\prime})$ and $(p_{\gamma+1}^{\prime},q_{\gamma+1})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime\prime})$. It follows that these objects are as required. Lastly, assume this strategy does not win, i.e. there is a game of length $\delta$. In this case, we claim that $\{(p_{\gamma}^{0},p_{\gamma}^{1})\;|\;\gamma\in\operatorname{Odd}\}$ is an antichain in $(\mathbb{P},b(R))^{2}$, obtaining a contradiction. Assume $(p^{0},p^{1})(b(R)^{2})(p_{\gamma}^{0},p_{\gamma}^{1}),(p_{\gamma^{\prime}}^{0},p_{\gamma^{\prime}}^{1})$ with $\gamma^{\prime}<\gamma$. Because $p^{0}b(R)p_{\gamma}^{0}$ and $p^{1}b(R)p_{\gamma}^{1}$, $(p^{0},q_{\gamma})R(p_{\gamma}^{0},q_{\gamma})$ and $(p^{1},q_{\gamma})R(p_{\gamma}^{1},q_{\gamma})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma^{\prime}}$ equally, but because $p^{0}b(R)p_{\gamma^{\prime}}^{0}$ and $p^{1}b(R)p_{\gamma^{\prime}}^{1}$, $(p^{0},q_{\gamma})R(p_{\gamma^{\prime}}^{0},q_{\gamma})R(p_{\gamma^{\prime}}^{0},q_{\gamma^{\prime}})$ and $(p^{1},q_{\gamma})R(p_{\gamma^{\prime}}^{1},q_{\gamma})R(p_{\gamma^{\prime}}^{1},q_{\gamma^{\prime}})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma^{\prime}}$ differently, a contradiction. ∎
We obtain an interesting statement regarding strongly distributive forcings (which contrasts the fact that consistently there can exist a ${<}\,\omega_{1}$-distributive, ccc. poset).
@@
    $\mathbb{P}$ is not strongly ${<}\,\delta$-distributive.
@@
Assume to the contrary that none of the above holds. As $\mathbb{P}$ is nontrivial, it adds some set $x$ of ordinals. Because $\mathbb{P}$ is in particular ${<}\,\delta$-distributive, $x\cap y\in V$ for every $y\in[x]^{<\delta}\cap V$. Thus $\mathbb{P}$ does not have the ${<}\,\delta$-approximation property. On the other hand, $\mathbb{P}$ preserves the $\delta$-cc. of itself by virtue of being strongly ${<}\,\delta$-distributive (by [4], Lemma 3.7), so $\mathbb{P}\times\mathbb{P}$ has the $\delta$-cc. and $\mathbb{P}$ has the ${<}\,\delta$-approximation property (by viewing $\mathbb{P}$ as an order on the product ×$\mathbb{P}\times\{0\}$), a contradiction. ∎
@@ 4. Ineffability Witnesses
    For all sufficiently large cardinals $\Theta$ and every $(\kappa,\lambda)$-list $f$, there is $M\prec H(\Theta)$ of size ${<}\,\kappa$ with $M\cap\kappa\in\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$.
@@
is stationary in $[\lambda]^{<\kappa}$. Using standard techniques we find $M\prec H(\Theta)$ with $\kappa,\lambda,f,b\in M$ such that $M\cap\kappa\in\kappa$ and $M\cap\lambda\in S$. Clearly $b\cap M=f(M\cap\lambda)$. Assume the submodel property holds and let $f$ be a $(\kappa,\lambda)$-list. Find $M\prec H(\Theta)$ of size $<\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$. We will show that $b$ is an ineffable branch for $f$. If $b$ is not an ineffable branch, by elementarity, there is $C\in M$ club in $[\lambda]^{<\kappa}$ such that for every $x\in C$, $f(x)\neq x\cap b$. $M$ contains a function $F$ such that every $x\in[\lambda]^{<\kappa}$ closed under $F$ is in $C$. Because $F\in M\prec H(\Theta)$, $M\cap\lambda$ is closed under $F$, so $M\cap\lambda\in C$ but $f(M\cap\lambda)=M\cap b$, a contradiction. ∎
Note that the previous Theorem works “list-by-list”, i.e. for any $(\kappa,\lambda)$-list $f$, $f$ has an ineffable branch if and only if there is $M\prec H(\Theta)$ of size ${<}\,\kappa$ with $M\cap\kappa\in\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$. We fix a definition corresponding to the previous Lemma:
@@
So in particular, we have shown that a $(\kappa,\lambda)$-list has an ineffable branch if and only if there is a $\lambda$-ineffability witness for $\kappa$ with respect to $f$. This enables us to show $\operatorname{ISP}(\delta,\kappa,\lambda)$ in the following way: Given a name $\dot{f}$ for a $\delta$-slender $(\kappa,\lambda)$-list, we construct (in the ground model) a related $(\kappa,\lambda)$-list $e$ (using the slenderness of $\dot{f}$) such that, given a $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ in the ground model, $M[G]$ is a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}^{G}$. Later, we will use witnesses with additional properties:
@@
    $M\cap\kappa\in\kappa$ is inaccessible and $[M\cap\lambda]^{<M\cap\kappa}\subseteq M$
@@
    $g(a)$ is a cofinal subset of $a\cap\kappa$ of ordertype ${<}\,a\cap\kappa$ if $a\cap\kappa$ is a singular ordinal.
@@
    $g(a)$ is an element of $[a]^{<a\cap\kappa}\smallsetminus F[a]$ if $a\cap\kappa$ is a regular ordinal and such an element exists.
@@
Since $\kappa$ is $\lambda$-ineffable, there is $M\prec H(\Theta)$ of size ${<}\,\kappa$ such that $\kappa,\lambda,f\in M$, $M\cap\kappa\in\kappa$ and for some $b\in M$, $M\cap b=f(M\cap\lambda)$. By the proof of Theorem 4.1 we can assume $x\subseteq M$ and that $M\cap\kappa$ is a strong limit. We will verify that only the last case can hold for $a:=M\cap\lambda$ (with the first case being excluded by assumption).
@@
    Assume $[a]^{<a\cap\kappa}\smallsetminus F[a]$ is nonempty. Then $g(a)$ is a subset of $M\cap\lambda$ of ordertype ${<}\,a\cap\kappa$. Proceed as before to show that $g(a)\in M$, which implies $g(a)\in F[a]$ by elementarity, a contradiction.
@@
We also need two results about the interaction of $\lambda$-ineffability witnesses and forcing: Given any set $M$ as well as any forcing order $\mathbb{P}$ and $\mathbb{P}$-generic filter $G$, let $M[G]$ consist of $\sigma_{G}$ for all $\sigma\in M$ that are $\mathbb{P}$-names. Note that if $\mathbb{P}\in M$ and $M\prec H(\Theta)$ for some large enough $\Theta$, $\Gamma_{\mathbb{P}}\in M$ and thus $G=\Gamma_{\mathbb{P}}^{G}\in M[G]$. The following is shown in [12], chapter III, theorem 2.11:
@@
Let $M\prec H(\Theta)$ of size ${<}\,\kappa$ with Mostowski-Collapse $\pi\colon M\to N$. Let $\mathbb{P}\in M$ be a forcing order and $G$ a $\mathbb{P}$-generic filter. The following are equivalent:
@@
Let $M\prec H(\Theta)$ of size ${<}\,\kappa$ with Mostowski-Collapse $\pi\colon M\to N$. Assume $\nu:=|M|=M\cap\kappa\in\kappa$ and $[M\cap\lambda]^{<\nu}\subseteq M$. Let $\mathbb{P}\in M$ be a poset of size $\lambda$ with the $\kappa$-cc.. Then the following holds:
@@
We will assume that $\mathbb{P}$ is a poset on $\lambda$. Let $\sigma\in M$ be such that $\sigma_{G}\in V$. By the forcing theorem and elementarity, there is $p\in G\cap M$ which forces $\sigma\in V$. Hence, again by elementarity, there exists $A\in M$ which is a maximal antichain in $\mathbb{P}$ such that for any $q\in A$ there is $y_{q}$ such that $q\Vdash\sigma=\check{y}_{q}$. Because $M\cap\kappa\in\kappa$ and $|A|<\kappa$ by the $\kappa$-cc. of $\mathbb{P}$, $A\subseteq M$ and thus $y_{q}\in M$ for every $q\in A$. Since $\sigma_{G}$ is one of the $y_{q}$, the statement follows. Now assume $x\in[\pi(\lambda)]^{\mu}\cap V[\pi[G\cap M]]$ for some $\mu<M\cap\kappa$. Let $\dot{x}$ be a $\pi(\mathbb{P})$-name for $x$ and for $\alpha\in\mu$, let $A_{\alpha}$ be a maximal antichain deciding the $\alpha$th value of $\dot{x}$. Then $|A_{\alpha}|<M\cap\kappa$ by the $M\cap\kappa$-cc. of $\pi(\mathbb{P})$. Let $f:\bigcup_{\alpha<\mu}\{\alpha\}\times A_{\alpha}\longrightarrow V$ be such that $f(\alpha,q)$ is the value $q$ decides for the $\alpha$th element of $\dot{x}$. Then (after some coding) $f\in[\pi(\lambda)]^{<M\cap\kappa}$, so $\pi^{-1}[f]\in M$ and $f=\pi(\pi^{-1}[f])\in N$. From $f$ and $\pi[G\cap M]$, we can recover $\dot{x}^{\pi[G\cap M]}$, so $\dot{x}^{\pi[G\cap M]}\in N[\pi[G\cap M]]$. ∎
The last result we need can be seen as dual to the lifting of elementary embeddings.
@@
Let $M\prec H(\Theta)$ with Mostowski-Collapse $\pi\colon M\to N$. Assume $\mathbb{P}\in M$ is a poset and $G$ is a $\mathbb{P}$-generic filter. The Mostowski-Collapse $\pi_{M[G]}$ is a function from $M[G]$ to $N[\pi[G\cap M]]$ and, if $M[G]\cap V=M$, it satisfies $\pi_{M[G]}\upharpoonright M=\pi$.
Proof.
We first show by induction on the rank of $\sigma\in M$ that $\pi_{M[G]}(\sigma^{G})=\pi(\sigma)^{\pi[G\cap M]}$. Assume $\sigma\in M$ is a $\mathbb{P}$-name and for every $(p,\tau)\in\sigma$ with $\tau\in M$, $\pi_{M[G]}(\tau^{G})=\pi(\tau)^{\pi[G\cap M]}$. Then
$\pi_{M[G]}(\sigma^{G})=\{\pi_{M[G]}(\tau^{G})\;|\;\tau^{G}\in\sigma^{G}\cap M[G]\}$
Let $\tau_{0}^{G}\in\sigma^{G}\cap M[G]$ and let $\tau_{1}\in M$ be such that $\tau_{1}^{G}=\tau_{0}^{G}$. Since $\tau_{1},\tau_{1}^{G},\sigma,\sigma^{G}$ and $G$ are in $M[G]$, there is $p\in G\cap M$ and $\tau_{2}\in M$ such that $\tau_{2}^{G}=\tau_{1}^{G}$ and $(p,\tau_{2})\in\sigma$. By assumption $\pi_{M[G]}(\tau_{2}^{G})=\pi(\tau_{2})^{\pi[G\cap M]}$. Furthermore, $\pi((p,\tau_{2}))=(\pi(p),\pi(\tau_{2}))\in\pi(\sigma)$ and $\pi(p)\in\pi[G\cap M]$, so $\pi(\tau_{2})^{\pi[G\cap M]}\in\pi(\sigma)^{\pi[G\cap M]}$. This shows $\pi_{M[G]}(\sigma^{G})\subseteq\pi(\sigma)^{\pi[G\cap M]}$. Let $\tau_{0}^{\pi[G\cap M]}\in\pi(\sigma)^{\pi[G\cap M]}$. Let $\pi(p)\in\pi[G\cap M]$ and $\pi(\tau_{1})$ be such that $\pi(\tau_{1})^{\pi[G\cap M]}=\tau_{0}^{\pi[G\cap M]}$ and $(\pi(p),\pi(\tau_{1}))\in\pi(\sigma)$. Thus $(p,\tau_{1})\in\sigma$ and $\tau_{1}^{G}\in\sigma^{G}$. Ergo $\tau_{0}^{\pi[G\cap M]}=\pi(\tau_{1})^{\pi[G\cap M]}=\pi_{M[G]}(\tau_{1}^{G})$, so $\pi(\sigma)^{\pi[G\cap M]}\subseteq\pi_{M[G]}(\sigma^{G})$. We also show $\pi_{M[G]}\upharpoonright M=\pi$ by induction on rank. Assume $\pi_{M[G]}(x)=\pi(x)$ for every $x\in y\cap M$ with $y\in M$ (note that $y\in M$ implies $y\cap M[G]=y\cap M$ since $y\subseteq V$). Then
$\pi_{M[G]}(y)=\{\pi_{M[G]}(x)\;|\;x\in y\cap M[G]\}=\{\pi(x)\;|\;x\in y\cap M\}=\pi(y)$
We now give a criterion for when a forcing order forces $\operatorname{ISP}$, generalizing the result and adapting the proof from [3]:

### v2

@@ 3. Strong Distributivity and Orders on Products
For better readability, we prove the result in a series of statements, showing where we apply which property. We let PP stand for the projection property, RP for the refinement property and MP for the mixing property. Let BS stand for basedness. We check two cases:
Assume there exist $q_{0},y_{0}$ such that $(p,q_{0})R(p,q)$ and $(p,q_{0})\Vdash\dot{x}=\check{y}_{0}$.
@@
$(MP)\;\;$ $\exists p_{0},p_{1},q^{\prime}(((p^{\prime},q^{\prime})R(p^{\prime},q))\wedge((p_{0},q^{\prime})R(p^{\prime},q_{0}))\wedge((p_{1},q^{\prime})R(p^{\prime},q_{1})))$
$p_{0},p_{1}(b(R))p^{\prime}$
$(BS)\;\;$ $(p^{\prime},q^{\prime})R(p^{\prime},q)R(p,q)$
@@
$((p_{0},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{0})\wedge((p_{1},q^{\prime\prime})\Vdash\dot{x}=\check{y}_{1})$
Assume case (1) does not hold, i.e. for all $q_{0},y_{0}$ with $(p,q_{0})R(p,q)$, $(p,q_{0})\not\Vdash\dot{x}=\check{y}_{0}$.
@@
$\exists(p_{0},q^{\prime}),y_{0}(((p_{0},q^{\prime})R(p,q))\wedge((p_{0},q^{\prime})\Vdash\dot{x}=\check{y}_{0}))$
@@
$p_{0},p_{1}(b(R))p$
$(p,q^{\prime\prime})R(p,q^{\prime\prime\prime})R(p,q)$
$(RP)\;\;$ $(p_{0},q^{\prime\prime})R(p_{0},q^{\prime\prime\prime})R(p_{0},q^{\prime})$
@@
Proof of Theorem 3.8.
Let $\dot{f}$ be a $(\mathbb{P}\times\mathbb{Q},R)$-name for a function such that some $(p,q)$ forces $\dot{f}\notin V$ and $\dot{f}\upharpoonright\check{u}\in V$ for every $u\in[V]^{<\delta}\cap V$. We will construct a winning strategy for INC in the completeness game of length $\delta$ played on $(\mathbb{Q},s(R,p))\upharpoonright q$. In any run $(q_{\gamma})_{\gamma\in\delta}$ of the game, we will construct $(p_{\gamma}^{0},p_{\gamma}^{1},y_{\gamma})_{\gamma\in\operatorname{Odd}\cap\delta}$ such that
@@
    $y_{\gamma}\in[V]^{<\delta}\cap V$ and the sequence $(y_{\gamma})_{\gamma\in\operatorname{Odd}\cap\delta}$ is $\subseteq$-increasing
@@
Assume the game has been played until some even ordinal $\gamma<\delta$. Let $y_{\gamma+1}^{\prime}:=\bigcup_{\alpha\in\gamma\cap\operatorname{Odd}}y_{\alpha}$, which has size $\delta$. $\dot{f}\upharpoonright y_{\gamma+1}^{\prime}$ is forced to be in $V$, so we can find $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime})R(p,q_{\gamma})$ which decides $\dot{f}\upharpoonright y_{\gamma+1}^{\prime}$. By the projection property, we can find $q_{\gamma+1}^{\prime\prime}$ such that $(p,q_{\gamma+1}^{\prime\prime})R(p,q_{\gamma})$ and $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime})$, so $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ also decides $\dot{f}\upharpoonright\check{y}_{\gamma+1}^{\prime}$. Because $\dot{f}$ is forced to be outside of $V$, there is $\beta$ such that $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ does not decide $\dot{f}(\check{\beta})$. Let $y_{\gamma+1}:=y_{\gamma+1}^{\prime}\cup\{\beta\}$. Then $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ does not decide $\dot{f}\upharpoonright\check{y}_{\gamma+1}$, so we find $q_{\gamma+1}^{\prime\prime\prime}$ and $p_{\gamma+1}^{0},p_{\gamma+1}^{1}$ such that $(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime\prime})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime})$ and $(p_{\gamma+1}^{0},q_{\gamma+1}^{\prime\prime\prime})$ and $(p_{\gamma+1}^{1},q_{\gamma+1}^{\prime\prime\prime})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma+1}$ differently. Lastly, use the projection property to obtain $q_{\gamma+1}$ such that $(p,q_{\gamma+1})R(p,q_{\gamma+1}^{\prime\prime})$ and $(p_{\gamma+1}^{\prime},q_{\gamma+1})R(p_{\gamma+1}^{\prime},q_{\gamma+1}^{\prime\prime\prime})$. It follows that these objects are as required.
Lastly, assume this strategy does not win, i.e. there is a game of length $\delta$. In this case, we claim that $\{(p_{\gamma}^{0},p_{\gamma}^{1})\;|\;\gamma\in\operatorname{Odd}\}$ is an antichain in $(\mathbb{P},b(R))^{2}$, obtaining a contradiction. Assume $(p^{0},p^{1})(b(R)^{2})(p_{\gamma}^{0},p_{\gamma}^{1}),(p_{\gamma^{\prime}}^{0},p_{\gamma^{\prime}}^{1})$ with $\gamma^{\prime}<\gamma$. Because $p^{0}(b(R))p_{\gamma}^{0}$ and $p^{1}(b(R))p_{\gamma}^{1}$, $(p^{0},q_{\gamma})R(p_{\gamma}^{0},q_{\gamma})$ and $(p^{1},q_{\gamma})R(p_{\gamma}^{1},q_{\gamma})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma^{\prime}}$ equally, but because $p^{0}(b(R))p_{\gamma^{\prime}}^{0}$ and $p^{1}(b(R))p_{\gamma^{\prime}}^{1}$, $(p^{0},q_{\gamma})R(p_{\gamma^{\prime}}^{0},q_{\gamma})R(p_{\gamma^{\prime}}^{0},q_{\gamma^{\prime}})$ (using the refinement property) and $(p^{1},q_{\gamma})R(p_{\gamma^{\prime}}^{1},q_{\gamma})$ $R(p_{\gamma^{\prime}}^{1},q_{\gamma^{\prime}})$ decide $\dot{f}\upharpoonright\check{y}_{\gamma^{\prime}}$ differently, a contradiction. ∎
We obtain an interesting statement regarding strongly distributive forcings (which contrasts the fact that consistently there can exist a $\omega_{1}-$distributive, ccc. poset).
@@
    $\mathbb{P}$ is not strongly $\delta-$distributive.
@@
Assume to the contrary that none of the above holds. As $\mathbb{P}$ is nontrivial, it adds some set $x$ of ordinals. Because $\mathbb{P}$ is in particular $\delta-$distributive, $x\cap y\in V$ for every $y\in[x]^{<\delta}\cap V$. Thus $\mathbb{P}$ does not have the $\delta-$approximation property. On the other hand, $\mathbb{P}$ preserves the $\delta$-cc. of itself by virtue of being strongly $\delta-$distributive (by [7], Lemma 3.7), so $\mathbb{P}\times\mathbb{P}$ has the $\delta$-cc. and $\mathbb{P}$ has the $\delta-$approximation property (by viewing $\mathbb{P}$ as an order on the product $\mathbb{P}\times\{0\}$), a contradiction. ∎
@@ 4. Ineffability Witnesses
    For all sufficiently large cardinals $\Theta$ and every $(\kappa,\lambda)$-list $f$, there is $M\prec H(\Theta)$ of size $\kappa$ with $M\cap\kappa\in\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$.
@@
is stationary in $\lambda]^{<\kappa}$. Using standard techniques we find $M\prec H(\Theta)$ with $\kappa,\lambda,f,b\in M$ such that $M\cap\kappa\in\kappa$ and $M\cap\lambda\in S$. Clearly $b\cap M=f(M\cap\lambda)$.
Assume the submodel property holds and let $f$ be a $(\kappa,\lambda)$-list. Find $M\prec H(\Theta)$ of size $\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$. We will show that $b$ is an ineffable branch for $f$. If $b$ is not an ineffable branch, by elementarity, there is $C\in M$ club in $\lambda]^{<\kappa}$ such that for every $x\in C$, $f(x)\neq x\cap b$. $M$ contains a function $F$ such that every $x\in[\lambda]^{<\kappa}$ closed under $F$ is in $C$. Because $F\in M\prec H(\Theta)$, $M\cap\lambda$ is closed under $F$, so $M\cap\lambda\in C$ but $f(M\cap\lambda)=M\cap b$, a contradiction. ∎
Note that the previous Theorem works “list-by-list”, i.e. for any $(\kappa,\lambda)$-list $f$, $f$ has an ineffable branch if and only if there is $M\prec H(\Theta)$ of size $\kappa$ with $M\cap\kappa\in\kappa$ such that $\kappa,\lambda,f\in M$ and for some $b\in M$, $b\cap M=f(M\cap\lambda)$.
We fix a definition corresponding to the previous Lemma:
@@
So in particular, we have shown that a $(\kappa,\lambda)$-list has an ineffable branch if and only if there is a $\lambda$-ineffability witness for $\kappa$ with respect to $f$. This enables us to show $\operatorname{ISP}(\delta,\kappa,\lambda)$ in the following way: Given a name $\dot{f}$ for a $\delta$-slender $(\kappa,\lambda)$-list, we construct (in the ground model) a related $(\kappa,\lambda)$-list $e$ (using the slenderness of $\dot{f}$) such that, given a $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ in the ground model, $M[G]$ is a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}^{G}$.
Later, we will use witnesses with additional properties:
@@
    $M\cap\kappa\in\kappa$ is inaccessible and $M\cap\lambda]^{<M\cap\kappa}\subseteq M$
@@
    $g(a)$ is a cofinal subset of $a\cap\kappa$ of ordertype $a\cap\kappa$ if $a\cap\kappa$ is a singular ordinal.
@@
    $g(a)$ is an element of $a]^{<a\cap\kappa}\smallsetminus F[a]$ if $a\cap\kappa$ is a regular ordinal and such an element exists.
@@
Since $\kappa$ is $\lambda$-ineffable, there is $M\prec H(\Theta)$ of size $\kappa$ such that $\kappa,\lambda,f\in M$, $M\cap\kappa\in\kappa$ and for some $b\in M$, $M\cap b=f(M\cap\lambda)$. By the proof of Theorem 4.1 we can assume $x\subseteq M$ and that $M\cap\kappa$ is a strong limit. We will verify that only the last case can hold for $a:=M\cap\lambda$ (with the first case being excluded by assumption).
@@
    Assume $a]^{<a\cap\kappa}\smallsetminus F[a]$ is nonempty. Then $g(a)$ is a subset of $M\cap\lambda$ of ordertype $a\cap\kappa$. Proceed as before to show that $g(a)\in M$, which implies $g(a)\in F[a]$ by elementarity, a contradiction.
@@
We also need two results about the interaction of $\lambda$-ineffability witnesses and forcing: Given any set $M$ as well as any forcing order $\mathbb{P}$ and $\mathbb{P}$-generic filter $G$, let $M[G]$ consist of $\sigma_{G}$ for all $\sigma\in M$ that are $\mathbb{P}$-names. Note that if $\mathbb{P}\in M$ and $M\prec H(\Theta)$ for some large enough $\Theta$, $\Gamma_{\mathbb{P}}\in M$ and thus $G=\Gamma_{\mathbb{P}}^{G}\in M[G]$.
The following is shown in [19], chapter III, theorem 2.11:
@@
Let $M\prec H(\Theta)$ with Mostowski-Collapse $\pi\colon M\to N$. Let $\mathbb{P}\in M$ be a forcing order and $G$ a $\mathbb{P}$-generic filter over $V$. The following are equivalent:
@@
Proof.
Let us sketch a proof.
Assume (1) holds. If $D\in M$ is open dense in $\mathbb{P}$, let $A\subseteq D$, $A\in M$, be a maximal antichain. Then $M$ contains the name $\tau:=\{(p,\check{p})\;|\;p\in A\}$. It follows that $p:=\tau^{G}\in M[G]\cap V=M$, which implies that $p\in A\cap M\cap G\subseteq D\cap M\cap G$.
Assume (2) holds. Whenever $D\in N$ is open dense in $\pi(\mathbb{P})$, $D=\pi(E)$ for some $E\in M$ which, by elementarity, is open dense in $\mathbb{P}$. Ergo $D\cap M\cap G$ contains some $p$. Then $\pi(p)\in\pi[M\cap G]\cap\pi(E)=\pi[M\cap G]\cap D$.
Assume (3) holds. Let $\tau$ be a $\mathbb{P}$-name for an element of $V$ lying in $M$. Then $M$ contains the open dense set $D$ of conditions $p$ forcing $\tau=\check{x}_{p}$ for some $x_{p}$. So $\pi(D)\in N$ is open dense in $\pi(\mathbb{P})$, which implies $\pi[M\cap G]\cap\pi(D)\neq\emptyset$ and thus that $D\cap M\cap G$ contains some $p$. The corresponding $x_{p}$ is in $M$ as well by elementarity, so $x_{p}=\tau^{G}\in M$ (since $p\in G$). ∎
@@
Let $M\prec H(\Theta)$ of size $\kappa$ with Mostowski-Collapse $\pi\colon M\to N$. Assume $\nu:=|M|=M\cap\kappa\in\kappa$ and $M\cap\lambda]^{<\nu}\subseteq M$. Let $\mathbb{P}\in M$ be a poset of size $\lambda$ with the $\kappa$-cc and let $G$ be $\mathbb{P}$-generic over $V$. Then the following holds:
@@
We will assume that $\mathbb{P}$ is a poset on $\lambda$.
Let $\sigma\in M$ be such that $\sigma_{G}\in V$. By the forcing theorem and elementarity, there is $p\in G\cap M$ which forces $\sigma\in V$. Hence, again by elementarity, there exists $A\in M$ which is a maximal antichain in $\mathbb{P}$ such that for any $q\in A$ there is $y_{q}$ such that $q\Vdash\sigma=\check{y}_{q}$. Because $M\cap\kappa\in\kappa$ and $|A|<\kappa$ by the $\kappa$-cc. of $\mathbb{P}$, $A\subseteq M$ and thus $y_{q}\in M$ for every $q\in A$. Since $\sigma_{G}$ is one of the $y_{q}$, the statement follows.
[…]
The last result we need can be seen as dual to the lifting of elementary embeddings. This is folklore, for a proof, see [8, Lemma 5.3.6].
@@
Let $M\prec H(\Theta)$ with Mostowski-Collapse $(N,\pi)$. Assume $\mathbb{P}\in M$ is a poset, $G$ a $\mathbb{P}$-generic filter and $M[G]\cap V=M$. Then the Mostowski-Collapse of $M[G]$ is given by $(N[\pi[G\cap M]],\pi_{M[G]})$, where $\pi_{M[G]}(\sigma^{G})=(\pi(\sigma))^{\pi[G\cap M]}$. Furthermore, $\pi_{M[G]}\upharpoonright M=\pi$.
We now give a criterion for when a forcing order forces $\operatorname{ISP}$, generalizing the result and adapting the proof from [6, Theorem 7.2]: